OpenStudy (anonymous):

Rewrite the equation 4x - 2y = 8 in slope-intercept form. -2y= -4x+ 8 2y= 4x+ 8 y= 2x- 4 y= 2x+ 4

OpenStudy (ajprincess):

slope-intercept form is y=mx+c canu do nw?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

After you @ajprincess

OpenStudy (anonymous):

do you just plug it in

OpenStudy (anonymous):

What are you looking for go ahead @ajprincess

OpenStudy (ajprincess):

To do that u need to bring y one side. U can do it by adding 2y both sides and subtractng 8 both sides. Then u get 4x-2y+2y-8=8+2y-8 4x-8=2y Nw divide by 2 both sides. then u get 2x-4=y Is that clear?

OpenStudy (anonymous):

thank you

OpenStudy (ajprincess):

yw.:)